2. In a hypothetical economy, farmer grows wheat, which he sells to a miller for Rs 40,000. The miller

turns the wheat into flour, which he sells to a baker for Rs 70,000. The baker turns the flour into

bread and sells to consumers for Rs 1,20,000. Consumers eat the bread.

a) Using Value added approach, compute GDP by mentioning value added at each activity level.

b) Who contributes more to GDP, farmer, miller or baker and how much?

c) Does this example suggest another way of computing GDP, other than Value added approach? If

yes, suggest the name of the approach, compute GDP and compare the result which you found

with the value of GDP computed by Value added approach.


1
Expert's answer
2021-12-08T19:24:16-0500

Solution:

a.). Derive value added at each stage:

A farmer sells to miller = 40,000 – 0 = 40,000

The miller sells to a baker = 70,000 – 40,000 = 30,000

The baker sells to consumers = 120,000 – 70,000 = 50,000

GDP (total valued added) = 40,000 + 30,000 + 50,000 = 120,000

 

b.). The baker contributes more to GDP than anyone else. The baker contributes 50,000 to GDP.

 

c.). Another approach of computing GDP is the expenditure approach.

Expenditure approach formula = C + I + G + X – M

Only C is available:

C = 120,000

GDP = C

GDP = 120,000

The value of GDP is the same compared to the value-added approach.